# Does synthetic data help low-resource Named Entity Recognition?

### v1

@@ Abstract
Named Entity Recognition(NER) for low-resource languages aims to produce robust systems for languages where there is limited labeled training data available, and has been an area of increasing interest within NLP. Data augmentation for increasing the amount of low-resource labeled data is a common practice. In this paper, we explore the role of synthetic data in the context of multilingual, low-resource NER, considering 11 languages from diverse language families. Our results suggest that synthetic data does in fact hold promise for low-resource language NER, though we see significant variation between languages.
Does synthetic data help low-resource Named Entity Recognition?
Gaurav Kamaththanks: work done during an internship at the National Research Council, Canada McGill University, Canada gaurav.kamath@mail.mcgill.ca                        Sowmya Vajjala National Research Council, Canada sowmya.vajjala@nrc-cnrc.gc.ca
@@ 1 Introduction
As a sequence labeling task, named entity recognition (NER) has historically one of the well-studied tasks in NLP. While much of the research in this direction assumes the presence of large datasets Sang and De Meulder (2003); Weischedel et al. (2011), there is also a large body of work focused on low-resource NER, i.e., developing methods for data scarce scenarios. Such scenarios can exist in two contexts: low-resource languages (e.g., Sohn et al., 2024) as well as low-resource domains within English (e.g., Santoso et al., 2024), both of which received considerable attention in NLP research in recent years. We approach the problem in the low-resource language context in this paper.
@@
Two common approaches to address such data scarce scenarios are: knowledge transfer and data augmentation. While knowledge transfer involves using cross-linguistically compatible text representations and/or transfer learning from a related high-resource setup, data augmentation approaches focus on increasing the amount of labeled data in the target setup. There are also large automatically constructed multilingual NER datasets that rely on large multilingual data sources such as Wikipedia Pan et al. (2017); Malmasi et al. (2022) and some such datasets have become a part of large multilingual benchmarks Asai et al. (2024). Recent research explored the usefulness of Large Language Models for such data augmentation and synthetic data creation in NLP Whitehouse et al. (2023); Li et al. (2023), and their use for NER is emerging Bogdanov et al. (2024).
In this background, we propose LLM based synthetic data generation using a small amount of gold examples as an alternative to relying on large scale, automatically labeled datasets for low-resource NER. By considering 11 languages from 3 language families- Danish, Swedish and Slovak from the Indo-European languages; Swahili, Kinyarwanda, Yoruba and Igbo from the Niger-Congo languages, and Kannada, Malayalam, Tamil, and Telugu from the Dravidian languages, we show that:
@@
    Creating synthetic data by prompting a large language model with a few manually labeled examples gives much better results than completely automatically labeled datasets.
@@
    Zero-shot transfer after training for NER on a related language is a surprisingly high baseline.
  3. 3.
    A small amount of high quality, manually annotated training data is often better than a large amount of synthetic data in low-resource NER performance.
To our knowledge, this is the first attempt to explore using large language models for synthetic data generation in low-resource NER, and the first to cover $>10$ languages. Based on our findings, we also summarize some best-practices for synthetic data generation with LLMs for low-resource NER.
We start with a review of related literature (Section 2) and describe our data generation approach in Section 3. The Results are detailed in Section 4 followed by further discussion on limitations and broader impact.
2 Related Work
Addressing NER for low resource scenarios has been one of the topics of interest in NLP for a long time. We can broadly identify three strands of research on this issue: One strand of research focused on how to achieve good transfer from a source language to the target language with a much smaller amount of training data. With the arrival of pre-trained language models, this extended into mono/multilingual NER models that could be used for zero/few-shot transfer on low(er) resource languages. Another direction of research focused on creating large scale datasets to address the data scarcity issue in low-resource languages. We summarize low-resource NER research along these two strands in this section.
Cross-lingual Transfer in NER:
Some of the early approaches in cross-lingual transfer relied on the use of dictionaries, knowledge bases, and parallel data to enrich the feature representations for NER in the target language Burkett et al. (2010); Darwish (2013). The use of non-lexical features such as POS tags, multilingual gazetteers Zirikly and Hagiwara (2015), phonological features Bharadwaj et al. (2016), translation and transliteration Huang et al. (2019); Johnson et al. (2019), and learning joint character representations between related languages Cotterell and Duh (2017) are some of the ideas explored in this context. Research advances in multilingual word and text embeddings, followed by pre-trained models, lead towards exploring the possibility of zero-shot or few-shot NER through a source NER model trained using labeled data from one or multiple languages Rahimi et al. (2019); Mueller et al. (2020); Zeng et al. (2022); Zhao et al. (2022); Yang et al. (2022); Zhou et al. (2022).
Creating Synthetic Datasets for Low-resource NER:
Parallel corpora or machine translated versions, and multilingual gazetteers were used as a means to create low-resource NER datasets from scratch by utilizing high quality NER datasets from a high resource language in past research Mayhew et al. (2017); Ni et al. (2017); Xie et al. (2018); Liu et al. (2021); Yang et al. (2022); Fetahu et al. (2022). WikiAnn Pan et al. (2017) is a large scale NER dataset for 282 languages, created by mining named entity data from Wikipedia’s cross-lingual links and has been a popular benchmark for cross-lingual and multilingual NER over the past years. However, some recent work Lignos et al. (2022) pointed to the low-quality example sets, and the lack of a human verified test set. Despite this, the dataset continues to be a common benchmark for cross-lingual NER Schmidt et al. (2022); Asai et al. (2024).
LLMs as Data Generators:
Some of the recent work explored using large language models such as GPT3.5 as data generators for NER. NuNER Bogdanov et al. (2024) uses GPT3.5 to automatically annotate a subset of the English part of C4, a cleaned common crawl dataset Raffel et al. (2020) and demonstrated its extendability to 7 other Indo-European languages. ProgGen Heng et al. (2024) proposed an approach to use GPT4 to generate domain relevant attributes which are then utilized to create domain specific training data and evaluated their approach on several English NER datasets across domains. Wang (2024).
Compared to these two approaches, we study synthetic data generation for NER across 11 languages covering three language families and five scripts, by taking high-quality labeled sentences as seed examples to generate new data, instead of relying on completely automatically created datasets. We also evaluate open LLMs along with GPT4 with this approach. Finally, we compare this approach with WikiAnn Pan et al. (2017); Rahimi et al. (2019), and with cross-lingual transfer from a related language with high-quality data.
3 Our Approach
We describe our general experimental setup for training and fine-tuning of NER models and the synthetic data generation process in this section.
3.1 Experimental Setup
At a high level, our approach to assess the potential of synthetic data for low-resource language NER involves the following steps:
  1. 1.
    Using the train split of some high quality, usually manually annotated NER dataset for a target language to generate synthetic data for that language with the help of an LLM, and
  2. 2.
@@ 3.1 Experimental Setup
    Comparing the performance of an NER model on the test split of the high quality dataset when trained on synthetic data from Step 1 and another model trained on the train split of the high quality dataset.
@@
For all of our experiments, we use the pre-trained version of XLM-RoBERTa-large Conneau et al. (2020) as our base model. We fine-tune this model on our synthetic and original training sets in two distinct settings.
@@
    In the second setting, we first fine-tune the model on the high quality NER data in a language related to the target language [^1] See Appendix for the full list of chosen related languages for all the target languages., so as to get a working NER model in the related language, and then further fine-tune this NER model on our synthetic or organic target language data.
@@
3.2 Implementation
3.3 Languages & Datasets
@@
We focus on 11 languages from diverse language families: Tamil, Kannada, Malayalam, Telugu (Dravidian), Kinyarwanda, Swahili, Igbo, Yoruba (Niger-Congo), Swedish, Danish and Slovak (Indo-European). Of these, Igbo, Yoruba, and Kinyarwanda are not a part of the 100 languages covered in the pre-trained XLM-Roberta model, but Swahili is a part of the pre-training languages.
We use the Universal NER dataset Mayhew et al. (2024) as our high quality, manually annotated dataset for Swedish, Danish and Slovak; MasakhaNER2 Adelani et al. (2022) for Kinyarwanda, Swahili, Igbo and Yoruba, and the Naamapadam dataset Mhaske et al. (2023) for Tamil, Kannada, Malayalam and Telugu. It is worth noting that while the first two datasets are completely manually annotated, the train and validation splits of the Naamapadam dataset are constructed using parallel corpora, and thus, in the authors’ admission, contain some amount of noise. Nevertheless, we choose it as our organic dataset, as (i) its test sets, which contain 500-1000 datapoints per language, are completely manually annotated, and (ii) it remains the largest NER resource for these four languages.
Crucially, all of these datasets cover largely identical NER categories, allowing for comparisons between them. The Universal NER and Naamapadam datasets cover persons, locations and organizations as categories; the MasakhaNER2 data covers these three categories, as well as dates.
3.4 Synthetic Data Generation Process
Our synthetic data generation process involves using large language models to generate new synthetic datapoints on the basis of existing, high quality NER annotations through the following steps:
  1. 1.
    First, we randomly sample $m$ data points from the train split of an organic NER dataset.
  2. 2.
  3. 3.
  4. 4.
Figure 1 shows a demonstration of this approach, with the exact prompt used. In our experiments, we set $m$ to 10, $n$ to 20, and $k$ to 500. This sets an upper cap of 5000 synthetic training data points, if every model response contains perfectly formatted data points.
Our method to extract correctly-formatted datapoints from raw model responses relies on regular expressions to identify potential JSON substrings in the model response; as shown in Table 1, model responses may contain data that is ill-formatted due to a number of reasons. We deem any data faulty for such reasons (among any other reasons, such as hallucinated labels) to be unusable, and do not include it in our synthetic train sets.
We experiment with three LLMs as our source of synthetic data. We use GPT-4222We use gpt-4-turbo, and all data generation with the model was conducted between November and December 2024. Achiam et al. (2023), which we assume to be the state of the art; we use Llama-3.1-8B-Instruct Dubey et al. (2024), as an instance of a much smaller, open-source instruction-tuned model; and finally, we use Cohere For AI’s aya-expanse-32b Dang et al. (2024), as an instance of a larger open source multilingual large language.
4 Results
4.1 Qualitative Analysis of LLM-Generated Data
Model Average % of Usable Train Datapoints
(Across All Languages)
GPT-4-turbo 97.0%
Llama-3.1-8B-Instruct* 59.3%
aya-expanse-32b 11.7%
Table 2: Average percentage of usable train datapoints produced by each model; we divide the number of usable synthetic datapoints produced for each language by 5000 (the synthetic dataset size if all responses were perfect), and then take the average across languages. *Llama-3.1’s rejected datapoints are often incomplete due to hitting new token limits, suggesting potentially higher capabilities under higher token limits.
(a) ner from scratch Setting
(b) ner fine-tuning Setting
Figure 2: NER model performance on a subset of languages when trained on increasingly large subsets of organic and synthetic data. aya-expanse-32b and Llama-3.1-8B-Instruct produced lower amounts of usable data; this is why they do not extend as far as the organic or GPT-4-produced data in fine-tuning data size. Performance at Fine-tuning Dataset Size = 0, only present in the ner fine-tuning setting, indicates zero-shot performance of a related-language NER model. See Appendix for similar plots of all 11 languages.
Table 1 illustrates the type of responses we get from the three models when we prompt them for synthetic data in the manner described in Section 3.4.
Broadly speaking, we find that GPT-4 provides the highest quality datapoints, followed by Llama-3.1-8B-Instruct, and lastly aya-expanse-32b. This is equally borne out by the proportion of usable training datapoints produced by each model, shown in Table 2.
While model responses from GPT-4 are almost always usable, we find a few persistent errors in responses from the other two models. We highlight some below:
  1. 1.
    Unequal token and tag lengths: Some responses from a model are largely well-formatted and coherent, but are rendered unusable because of a mismatch in their numbers of tokens and tags.
  2. 2.
    Run-on responses and incomplete data: Some responses from the models result in endless repeated sequences; in other cases, responses hit new token limits before arriving at complete datapoints.
  3. 3.
    Empty responses and prompt continuations: These were specific to aya-expanse-32b; model responses to the generation prompt often either terminated with <EOS> tokens, or simply continued the generation prompt, as shown in the example in Table 1.
Given the wide range of languages we test, and the cost and difficulty of finding trained annotators to manually annotate samples from all 11 languages, we do not source human annotations of the synthetic data to compare against. Instead, as detailed in Section 3.1, we assess the overall quality and viability of this synthetic data by measuring the performance of an NER model on a high quality, manually-annotated test set, when trained on the synthetic data.
4.2 Training on Synthetic Data
Figure 2 shows our results on a subset of languages—Danish, Tamil and Igbo—when using synthetic data from different models, in both the ner from scratch and ner fine-tuning settings (see Appendix for such plots across all languages). As the figure illustrates, in all cases, models trained on LLM-generated data struggle to compete with those trained on the original, organic train sets. Perhaps more surprisingly, in the ner fine-tuning setting, zero-shot transfer onto the target language outperforms the same models after they have been further fine-tuned on synthetic target language data, suggesting that in such cases, the synthetic data is actually detrimental to their performance.
In terms of comparisons between data-generating LLMs, these results also confirm our initial qualitative observations (see Section 4.1). Models trained on synthetic data from GPT-4 do better than those trained on synthetic data from Llama-3.1-8B-Instruct, which in turn outperform the aya-expanse-32b equivalents—suggesting that the quality of data produced by the three LLMs follows the same order.
4.3 WikiAnn
Language n.f.s. F1 n.f.t. F1 Data Size
Kannada WikiANN 4.5e-3 0.77 100
GPT-4 0.26 0.65 100
GPT-4 0.65 0.68 4861
Naamapadam 0.47 0.79 100
Naamapadam 0.76 0.79 5000
Danish WikiANN 0.72 0.71 20000
GPT-4 0.60 0.68 4857
Universal NER 0.83 0.85 4383
Yoruba WikiANN 0.07 0.21 100
GPT-4 0.26 0.43 100
GPT-4 0.53 0.56 4761
MasakhaNER 2 0.20 0.50 100
MasakhaNER 2 0.79 0.82 5000
Swahili WikiANN 0.50 0.59 1000
GPT-4 0.74 0.78 1000
GPT-4 0.75 0.79 4900
MasakhaNER 2 0.69 0.85 1000
MasakhaNER 2 0.92 0.90 5000
Table 3: Performance of NER models trained on WikiANN, synthetic data from GPT-4, and high quality ‘organic’ data, for a handful of languages. See Appendix for full table of results. n.f.s: ner from scratch setting; n.f.t: ner fine-tuning setting.
As a further evaluation of the viability of synthetically produced data for lower-resource language NER, we also extend our comparison to the WikiANN dataset. The WikiANN dataset is a massively multilingual NER benchmark, comprising data from 176 languages Pan et al. (2017); Rahimi et al. (2019) [^3] As Lignos et al. (2022) also note, strictly speaking, the original version of WikiANN put together by Pan et al. (2017) contains data from 282 languages; the version of the dataset commonly downloaded from Huggingface, however, and put together by Rahimi et al. (2019), contains data from 176 languages. In this work, we refer to the latter when referring to the WikiANN dataset.. Though it is often used as a multilingual NER benchmark (e.g. Asai et al., 2024), however, the WikiANN dataset in fact does not include manual annotations, and is instead produced by scraping Wikipedia data. Consequently, the dataset represents a different sort of synthetic data—one generated not from LLMs, but instead from scraping knowledge bases.
Table 3 shows the performance of NER models trained on WikiANN data, as compared with those trained on our organic and LLM-produced synthetic data. The sizes of the WikiANN train sets vary significantly between different languages, meaning we often cannot assess the quality of the data in the context of training sets containing over 1000 datapoints (e.g. Kannada and Yoruba, whose WikiANN train sets contain only 100 datapoints). In such cases, however, we compare model performance when trained on equally small amounts of organic or LLM-produced synthetic data.
As Table 3 indicates, training on WikiANN data yields NER models that perform considerably worse than those trained on organic data, and often worse than those trained on synthetic data produced by GPT-4 (when data size is equal between the two). The WikiANN data for some languages appears completely ill-suited for training an NER model in the ner from scratch setting, while, perhaps unsurprisingly, organic data produces the best results.
5 Conclusions and Discussion
Our results suggest three broad findings around the utility of LLM-generated synthetic data for low resource language NER.
  1. 1.
    The first, and perhaps most significant, is that a small amount of carefully annotated data yields better performance than a large amount of synthetic data. As is evident in Figure 2, even 100 manually annotated datapoints can yield NER models that cannot be matched by models trained on much larger amounts of synthetic data.
  2. 2.
    The second is that, in many cases, zero-shot transfer from a related-language NER model is a high baseline, and that further training such a model on synthetic data may actually lower its performance.We find this to be true in the case of all languages tested except the four Niger-Congo languages—Igbo, Yoruba, Swahili and Kinyarwanda. In these cases, it is worth noting that the overall baselines are lower, presumably because these languages are all lower resource than the others tested.This may explain why synthetic data yields performance gains over the zero-shot baseline, though it does not change the trend of a small amount of manually annotated data yielding far better performance.
  3. 3.
    Lastly, we find that LLM-generated synthetic data often produces better results than automatically labelled data. Despite the fact that it falls short of manually annotated data, LLM-generated data often still yields better model performance than WikiANN data, as shown in the case of Kannada, Yoruba and Swahili. This echoes the findings by Lignos et al. (2022), who arrive at similarly negative findings around the data quality of WikiANN; consequently, they further reinforce the need for manually annotated gold test sets in benchmarking NER for lower resource languages.
6 Limitations
Although we experimented with many languages, the nature of the NER datasets used is relatively simple. Thus, we don’t know if the general conclusions, especially about the quality of synthetic data, will extend to scenarios where there are many entity categories. While we did study datasets covering more than one language family, the selection of language is far from extensive, and is also constrained by the availability of human labeled test data. Finally, to keep the experiments under control, we explored a limited set of methods for fine-tuning and synthetic data generation. So, our findings should be viewed after taking these aspects into consideration.
7 Ethics and Broader Impact
We used publicly available datasets with human-annotated and automatically labeled data, and also created synthetically generated datasets as a part of this work. The models built using such artificially created datasets should always be validated with a human-labeled data. We did not involve any human participants in this study. All the code and generated datasets will be provided as supplementary material to support reproducible research.
Acknowledgments
This research was conducted at the National Research Council of Canada, thereby establishing a copyright belonging to the Crown in Right of Canada, that is, to the Government of Canada.
Appendix A Related-Language Model Details

### v2

# Does Synthetic Data Help Named Entity Recognition for Low-Resource Languages?

@@ Abstract
In this paper, we explore whether synthetic datasets generated by large language models are useful for low-resource named entity recognition, considering 11 languages from diverse language families. Our results suggest that synthetic data created with seed human labeled data is a reasonable choice when there is no available labeled data, and is better than using automatically labeled data. However, a small amount of high-quality data, coupled with cross-lingual transfer from a related language, always offers better performance.
Does Synthetic Data Help
Named Entity Recognition for Low-Resource Languages?
Gaurav Kamaththanks: Work done during an internship at the National Research Council, Canada.1, 2  Sowmya Vajjala3 1McGill University, Canada  2Mila - Quebec AI Institute, Canada 3National Research Council, Canada gaurav.kamath@mail.mcgill.casowmya.vajjala@nrc-cnrc.gc.ca
@@ 1 Introduction
Named Entity Recognition (NER) for low-resource languages aims to produce robust systems for languages with limited labeled training data available, and has been an area of increasing interest within natural language processing (NLP). Two common approaches to address this data scarcity are cross-lingual transfer and data augmentation/synthesis; recent research has in particular explored the usefulness of large language models (LLMs) for such data augmentation and synthetic data creation in NLP Whitehouse et al. (2023); Li et al. (2023), while their use for NER is also emerging Bogdanov et al. (2024).
In this background, we propose LLM-based synthetic data generation using a small amount of gold examples (Figure 1) as an alternative to relying on automatically created datasets for low-resource NER. With experiments covering 11 languages, we show that
  1. 1.
    Even a small amount of human annotated data can yield far better performance than much larger amounts of synthetic data.
  2. 2.
    Zero-shot transfer from a related language can provide high baselines for low-resource language NER.
  3. 3.
    Synthetic data generated by prompting an LLM with a few high quality examples (Figure 1) could be better than using automatically labeled datasets when training low-resource NER models.
@@
We start with a review of related literature (Section 2) and describe our data generation approach and experimental setup in Section 3, followed by a discussion of the results (Section 4), limitations (Section 6) and broader impact (Section 7).
2 Related Work
NER in low resource settings has long been a topic of interest in NLP. Significant research examines cross-lingual transfer from a high resource source language to a lower-resource target language for the task Rahimi et al. (2019); Mueller et al. (2020); Zeng et al. (2022); Zhao et al. (2022); Yang et al. (2022); Zhou et al. (2022), while other approaches have explored the creation of synthetic datasets through e.g. parallel corpora or machine translation Mayhew et al. (2017); Ni et al. (2017); Pan et al. (2017); Xie et al. (2018); Liu et al. (2021); Yang et al. (2022); Fetahu et al. (2022).
More recent work has explored using LLMs such as GPT-3.5 and GPT-4 as data generators for NER Bogdanov et al. (2024); Heng et al. (2024). We build on such work, but differ from their methods. Our data generation process uses high quality examples as seeds, and we not only evaluate different LLMs (both open and closed-source), but also experiment with 11 languages covering three language families and five base scripts.
3 Our Approach
At a high level, our approach involves two steps:
@@ 3 Our Approach
    Using the train split of a high quality (usually manually annotated) NER dataset for a target language to generate synthetic data for that language with the help of an LLM (Section 3.1); and then
@@
    Comparing the performance of an NER model on the test split of the high quality dataset when trained on synthetic data from Step 1 and another model trained on the train split of the same high quality dataset (Section 3.2).
3.1 Synthetic Data Generation:
Our synthetic data generation process (shown in Figure 1) involves using LLMs to generate new synthetic data points on the basis of existing, high quality NER annotations as described below:
  • First, we randomly sample $m$ data points from the train split of an organic (i.e. non-synthetic) NER dataset.
In our experiments, we set $m$ to 10, $n$ to 20, and $k$ to 500. This sets an upper cap of 5000 synthetic training data points, if every model response contains perfectly formatted data points. We present and solicit data structured as JSON strings to the LLMs, and extract well-formatted samples from model responses using regular expressions. Appendix A provides further details about this process.
We compare three LLMs as our source of synthetic data: GPT-4111We use gpt-4-turbo, and all data generation with the model was conducted between September and December 2024. Achiam et al. (2023), which we assume to be the state of the art; Llama-3.1-8B-Instruct Dubey et al. (2024), as a much smaller, open-source instruction-tuned model; and finally, aya-expanse-32b Dang et al. (2024), as a larger open source multilingual LLM.
3.2 Training NER models:
@@
For all experiments, we use the pre-trained version of XLM-RoBERTa-large Conneau et al. (2020) as our base model and fine-tune it on our synthetic and organic training sets in two distinct settings.
@@
    In the second setting, we first fine-tune the model on the high quality NER data in a language related source language [^2] See Table 2 in Appendix B for the full list of chosen related languages for all the target languages., and then further fine-tune this NER model on our synthetic or organic target language data.
@@
Languages & Datasets:
We focus on 11 languages from diverse language families: Tamil, Kannada, Malayalam, Telugu (Dravidian), Kinyarwanda, Swahili, Igbo, Yoruba (Niger-Congo), Swedish, Danish and Slovak (Indo-European). Of these, Igbo, Yoruba, and Kinyarwanda are not among the 100 languages in the XLM-Roberta pre-training corpus. We use the Universal NER dataset Mayhew et al. (2024) as our high quality, manually annotated dataset for Swedish, Danish and Slovak; MasakhaNER2 Adelani et al. (2022) for Kinyarwanda, Swahili, Igbo and Yoruba; and the Naamapadam dataset Mhaske et al. (2023) for Tamil, Kannada, Malayalam and Telugu.
While the first two datasets are completely manually annotated, the train and validation splits of the Naamapadam dataset are constructed using parallel corpora, and thus contain some noise. Nevertheless, we choose it as our organic dataset, as (i) its test sets, which contain 500-1000 datapoints per language, are completely manually annotated, and (ii) it remains the largest NER resource for these four languages. Crucially, all of these datasets cover largely identical NER categories, allowing for comparisons between them.
Additionally, we compare models trained on LLM-generated data with those trained using WikiANN Pan et al. (2017); Rahimi et al. (2019), a large, automatically created NER dataset based on Wikipedia cross-linking, as it covers the 11 languages we study. This dataset represents a different form of synthetic data—one generated not from LLMs, but instead from scraping knowledge bases. Although the dataset has no manual annotations, it is frequently used as a standard low-resource NER benchmark Schmidt et al. (2022); Asai et al. (2024).
4 Results
4.1 Synthetic Data Generation
We generate the synthetic datasets following the process described in Section 3.1. While model responses from GPT-4 are almost always usable, we found recurring errors in responses from the other two models. Some of these errors are described in Table 1 in Appendix A; we discard such instances when compiling our synthetic datasets from model responses. The average percentage of usable training datapoints from GPT-4, Llama-3.1 and aya-expanse are 97%, 59.3% and 11.7% respectively. [^3] Llama-3.1’s rejected datapoints are often incomplete due to hitting new token limits, suggesting potentially higher capabilities under higher token limits. We assess the overall quality and viability of this synthetic data by measuring the performance of an NER model on a high quality, manually-annotated test set, when trained on the synthetic data.
4.2 Training on Synthetic Data
Figure 2 shows our results when using synthetic data from different models, in both the ner from scratch and ner fine-tuning settings. While the models trained on organic data in the ner from scratch setting always perform better than synthetic data based models, we find that models trained on GPT-4-generated data come the closest to models trained on organic data. We also find that more synthetic data is not necessarily useful; for some languages, we see a saturation after about 1000 data points, and for some, we also notice a drop in performance with more data.
Perhaps more surprisingly, in the ner fine-tuning setting, we notice that zero-shot transfer from a related language outperforms the same models after they have been further fine-tuned on synthetic target language data. This suggests that in some cases where an NER model for a related language exists, synthetic data in target languages may actually be detrimental to overall performance.
(a) ner from scratch Setting
(b) ner fine-tuning Setting
Figure 2: NER model performance when trained on increasingly large subsets of training data. aya-expanse-32b and Llama-3.1-8B-Instruct produced lower amounts of usable data; this is why they do not extend as far as organic or GPT-4-produced data in fine-tuning data size. Performance at Fine-tuning Dataset Size = 0, only present in the ner fine-tuning setting, indicates zero-shot performance of a related-language NER model.
Comparison with WikiAnn:
In most cases, when dataset size is comparable, training on WikiANN data in the ner from scratch setting yielded NER models that perform considerably worse than those trained on synthetic data from GPT-4. For the four Niger-Congo languages, GPT-4 generated data gave superior results even in the ner fine-tuning setting (see Table 3 in Appendix C for the detailed results).
5 Conclusions and Discussion
Our results lead us to three main conclusions around the utility of LLM-generated synthetic data for low resource language NER.
  1. 1.
    A small amount of carefully annotated data yields better performance than a large amount of synthetic data. As is evident in Figure 2, even 100 manually annotated data points can yield NER models that cannot be matched by models trained on much larger amounts of synthetic data.
  2. 2.
    In many cases, zero-shot transfer from a related-language NER model is a high baseline, and that further training such a model on synthetic data may even lower the performance.
  3. 3.
    Despite the fact that it falls short of manually annotated data, LLM-generated data often still yields better model performance than WikiANN, especially for the more low-resource languages among the ones we studied. This echoes the findings by Lignos et al. (2022), who arrive at similarly negative findings around the data quality of WikiANN.
Overall, while showing how synthetic data from LLMs can help train NER models from scratch for low resource languages, our results reinforce the need for manually annotated gold test sets in benchmarking NER for lower resource languages.
6 Limitations
Although we experimented with many languages, the nature of the NER datasets used is relatively simple, containing only three or four entity categories (persons, locations, organizations and dates). Thus, we don’t know if the general conclusions, especially about the quality of synthetic data, will extend to scenarios where there are many entity categories. While we did study datasets covering more than one language family, the selection of language is far from extensive, and is also constrained by the availability of human labeled test data. Finally, to keep the experiments under control, we explored a limited set of methods for fine-tuning and synthetic data generation. Our findings should be viewed after taking these aspects into consideration.
7 Ethics and Broader Impact
We used publicly available datasets with human-annotated and automatically labeled data, and also created synthetically generated datasets as a part of this work. The models built using such artificially created datasets should always be validated with a human-labeled data. We did not involve any human participants in this study. All the code and generated datasets is provided at this github repository to support reproducible research: https://anonymous.4open.science/r/low-resource-syn-ner-A1C7/.
Appendix A Synthetic Data Generation
@@
{m examples}
Table 1 shows some of the examples of the different types of responses to these prompts.
Appendix B Related-Language Model Details